A woman rides a carnival Ferris wheel at radius 18 m, completing 5.8 turns about its horizontal axis every minute. What are (a) the period of the motion, and the magnitude of her centripetal acceleration at (b) the highest point and (c) the lowest point?
Answer #1

a)

T = 60 / 5.8 = 10.3448 s

b)

a = r w^2 = 18* ( 2* 3.14 / 10.3478)^2

a = 6.63 m/s^2

c)

a = rw^2 = 6.63 m/s

=======

Comment before rate in case any doubt, will reply for sure.. goodluck